# Unveiling Molecular Moieties through Hierarchical Graph Explainability

### v2

@@ Abstract
Background: Graph Neural Networks (GNN) have emerged in very recent years as a powerful tool for supporting in silico Virtual Screening. In this work we present a GNN which uses Graph Convolutional (GC) architectures to achieve very accurate multi-target screening. We also devised a hierarchical Explainable Artificial Intelligence (XAI) technique to catch information directly at atom, ring, and whole molecule level by leveraging the message passing mechanism. In this way, we find the most relevant moieties involved in bioactivity prediction.
Results: We report a state-of-the-art GNN classifier on twenty Cyclin-dependent Kinase targets in support of VS. Our classifier outperforms previous SOTA approaches proposed by the authors. Moreover, a CDK1-only high-sensitivity version of the GNN has been designed to use our explainer in order to avoid the inherent bias of multi-class models. The hierarchical explainer has been validated by an expert chemist on 19 approved drugs on CDK1. Our explainer provided information in accordance to the docking analysis for 17 out of the 19 test drugs.
Conclusion: Our approach is a valid support for shortening both the screening and the hit-to-lead phase. Detailed knowledge about the molecular substructures that play a role in the inhibitory action, can help the computational chemist to gain insights into the pharmacophoric function of the molecule also for repurposing tasks.
footnotetext: $*{}^{*}$corrisponding author salvatore.contino01@unipa.it (Salvatore Contino), uperricone@fondazionerimed.com (Ugo Perricone)footnotetext: ${}^{\ddagger}$hese authors contributes equally.footnotetext: ${}^{1}$epartment of Engineering, University of Palermo, 90128 Palermo, Italyfootnotetext: ${}^{2}$olecular Informatics Group, Fondazione Ri.MED, 90128 Palermo, Italy
@@ 1 Background
[…]
@@
In this context, Graph Neural Networks (GNN) [18, 12] have emerged as powerful tools in Drug Discovery and Virtual Screening due to their ability to model complex molecular structures and interactions. By leveraging the graph-based representation of molecules to capture local and global structural information, GNNs offer a promising approach for predicting molecular properties, bioactivity and drug-target interactions [9, 28, 24, 26, 20, 5, 23].
[…]
@@
In general, Grad-CAM is applied only to the last layer of the classifier whose feature maps convey aggregated position-independent information on the features extracted by the deep layers of the network. Rather than providing a global explanation of the classifier behaviour, the key idea behind our approach is trying to explain how the different molecular moieties are involved in predicting a high degree bioactivity. These structures are expected to play a crucial role in binding to the target. As a consequence, we used a hierarchy of explainers placed at different layers to catch information directly at atom, ring, and whole molecule level by leveraging the message passing mechanism that is used in GNNs. The explainer hierarchy for a target is applied purposely to a binary version of our classifier. Such a network has a similar layer arrangement as the 20-target network, but the layers’ sizes have been designed ad hoc, and a novel training on the selected target has been performed to achieve high accuracy, while avoiding the inherent bias of a multi-target prediction with respect to the different classes. In this paper we present the implementation of the whole explainability procedure for CDK1, which is a well-known regulator of the cell cycle, and has a crucial role in tumor genesis. Our explainability results for nineteen well known drugs that are active on CDK1 where manually validated by expert computational chemists, and this was a very time consuming task. The proposed approach is exactly the same for each target so this choice does not reduce the generality of our results.
@@ 2.1 Data preparation
The set of atom features was later augmented with the 3D coordinates of each atom in active configuration, for the sake of training the single-class architecture used for the hierarchical explainability analysis. These coordinates were calculated by maximising the local minimum of the active conformation for each individual molecule. Edges information concerns the type of bond (double or triple bond) and whether the bond is within a cycle. The size of each molecular graph had several nodes and arcs which vary according to the size of the molecule.
[…]
The final dataset consists of 89373 small molecules and has a 1:100 active/inactive ratio for the smallest class that is CLK2. This ratio allows profound assessing of the model generalization abilities as it has more or less the same size of a real world screening task. Training, validation, and test sets were created using a 80%:10%:10% ratio for each target, in order to minimize the comparison bias.
@@ 2.2 Graph Convolutional Neural Network
The single target architecture differs significantly from the first one as regards both the number and the size of each layer. It is well known that the global efficacy of a model for Virtual Screening is measured through the ability of correctly prioritizing the True Positives, so the main aim of the network is providing a very high sensitivity with respect to the inhibitors of the target that is CDK1 in our experiments. In turn, stressing the network towards high sensitivity performance, makes it able to provide precise explanations about the molecular moieties involved in the inhibition. To achieve this result we operated in two ways: enrichment of the information encoded at each node, and accurate design of the model capacity through a change in the dimensions of the layers. The 3-D coordinates of each atom were computed by maximising the local minimum of the active conformation of the molecule, and they were added to the node features. Moreover, we selected a tighter network than the multi-target one because the task is only a binary classification.
The GC layers in the single target network elaborate the features in a U-Net [17] fashion, first enlarging, and then tightening the channels at each layer until a final 32 features representation is obtained. Seven GC blocks are used where the channels of each GC layer are arranged as follows: 32, 64, 128, 256, 128, 64, and again 32. Layer normalization and ReLU are obviously used in each block. After the GAP, four fully connected layers where used whose channels are respectively 2, 32, 16, and finally 1 that is fed to a sigmoid to estimate the probability for the input to be active on the target. The whole architecture of the single target network is reported in Figure 2 along with an example of the output of the explainers used in the proposed hierarchy.
@@ 2.3 Explainability analysis
Bioactivity prediction is a classification task where molecules are labelled as positive or negative if they trigger “globally” a particular chemical event or not. Many factors are involved in bioactivity apart from the presence of the classical pharmacophoric structures: molecular chirality, hydrosolubility, use of conformational restriction in the drug design, and so on. All of them can be captured by training the GCNN on a three-dimensional molecular representation with rich information at the nodes, but can not be directly related to just “positive” or “negative” Grad-CAM values at each node. As a consequence we removed ReLU for our explainability task. Moreover, pharmacophoric information is strictly related to the presence in the molecule of particular structures that span atoms’ neighborhoods with different size. In view of this consideration we evaluated Grad-CAM multiple times at different depths in the convolutional segment of the network. We used the classical Grad-CAM arrangement after the last layer of the network that is the 7-th one. This explainer is aimed at highlighting relevant features globally, and it is not able to provide information about subtle moieties like acceptors or donors. To reach this aim, we placed also a Grad-CAM explainer at the very first GC layer to consider the atoms’ impact in a 1-neighbourhood that is at a bond level. Another explainer was placed after the third layer to catch information related to 3-neighborhoods. At this level, as we analyze the 3-neighborhood of each atom, the action of an entire substructure like aromatic ring or carbonyl ring can be detected.
@@
The resulting $L_{Grad-CAM}$ distribution has been min-max scaled, and only contributions higher than $0.7$ have been regarded as relevant. The RDKithttps://www.rdkit.org/ software package has been used to devise pharmacophoric labels for each molecular structure, and the final map reports only relevant contributions colored accordingly to the labels. In particular, H-bond acceptor or donor groups are highlighted in yellow, hydrophobic groups are colored in blue, and the aromatic groups are green. Lastly, atoms without an explicit pharmacophoric label where $L_{Grad-CAM}>0.7$ are labelled as “relevant”, and are colored in red. In Section LABEL:exp-results a detailed analysis is reported about the structures highlighted by our hierarchical Grad-CAM explainer.
@@ 3.1 Multi-class multi-target classification
Results for the 20-target classifier are shown in Table 1 and Table 2. The first table provides the performance metrics of our network as a classifier. The Balanced Accuracy assesses the network’s ability yo make correct predictions despite the class imbalance, while the Sensitivity (also named Recall) focuses solely on the network’s performance in correctly classifying active molecules. The F1-score quantifies how well the model can correctly identify positive instances (Recall) while avoiding false positives (Precision). On the other hand, the Area Under the Curve (AUC) provides a global measure of the model’s ability to discriminate between positive and negative instances, and it is computed from the curve that plots the Sensitivity vs the False Positive Rate varying the threshold to discriminate positives and negatives.
@@
In our results, both the accuracy and sensitivity values indicate that our multi-target network identifies correctly both the positive and negative samples. Particularly, sensitivity increased significantly with respect to the EMBER classifier where it ranged between 0.22 and 0.84. Now, sensitivity has an average value of 0.904, showing that it is much easier for the network to classify positive samples. The comparison between the two networks is summarised in Figure 3 showing the sensitivity for all targets. It is notable that there was a huge improvement in the performance related to sensitivity. Particularly, sensitivty on the CLK2 target raised from 0.22 to 0.938, as it is shown also in Table 1.
Figure 3: Sensitiviy comparison for all the targets between the EMBER and our GC multi-target classifier.
F1-Score values show that the network is able to overcome the dataset imbalance while simultaneously reducing the number of false negatives and false positives. The use of a properly balanced loss made by a weighted cross entropy and a soft F-score worked fine for the classification task. Finally, also the values for TP/P and EF are similar to the ones reported for the EMBER classifier.

### v3

@@ Abstract
Background: Graph Neural Networks (GNN) have emerged in very recent years as a powerful tool for supporting in silico Virtual Screening (VS). In this work we present a GNN which uses Graph Convolutional (GC) architectures to achieve very accurate multi-target screening. As novelty compared to state-of-the-art we devised a hierarchical eXplainable Artificial Intelligence (XAI) technique to catch information directly at atom, ring, and whole molecule level by leveraging the message passing mechanism. In this way, we find the most relevant moieties involved in bioactivity prediction.
Results: We report a state-of-the-art GNN classifier on twenty Cyclin-dependent Kinase targets in support of VS. Our classifier outperforms previous SOTA approaches proposed by the authors. Moreover, a CDK1-only high-sensitivity version of the GNN has been designed to use our explainer in order to avoid the inherent bias of multi-class models. The hierarchical explainer has been validated by comparison with literature, experimental and docking data on 19 approved drugs on CDK1. Our explainer provided information in accordance to crystal structures and the docking results analysis for 17 out of the 19 test drugs.
Conclusion: Our approach is a valid support for shortening both the screening and the hit-to-lead phase. Detailed knowledge about the molecular substructures that play a role in the inhibitory action, can help the computational chemist to gain insights into the pharmacophoric function of the molecule also for repurposing tasks. Scientific Contribution Statement The core scientific innovation of our work is the use of a hierarchical XAI approach on a GNN trained for a ligand-based VS task. The application of the hierarchical explainer allows for eliciting also structural information. Apart from the novelty of using a hierarchy of explainers that take into account wider and wider portions of the molecule, the other computational VS methods are not able to achieve our result, which allow for revealing pharmacophoric chemical moieties (acceptor/donor H-bond, aromatic, hydrophobic), while also integrating the contributions of moieties that might be relevant in the stabilisation phenomena of the interaction with the protein target even if they are not directly involved with pharmacophoric activity.
footnotetext: corrisponding author salvatore.contino01@unipa.it (Salvatore Contino)footnotetext: These authors contributes equally.footnotetext: 1Department of Engineering, University of Palermo, 90128 Palermo, Italyfootnotetext: 2Molecular Informatics Group,Fondazione Ri.MED, 90128 Palermo, Italy
@@ 1 Background
The development of a new drug is a long and time-consuming process that, despite advances in technology and computer applications, is fragmented into several steps before being able to identify a lead compound that can be tested on in vivo models. Over the last decade, the use of deep learning has provided support to the Virtual Screening (VS) campaigns for the efficient selection of hit compounds by strengthening predictions through the use of massive data in the training model phase [12, 4, 26, 3, 35]. Indeed, being able to correctly prioritise active molecules during the screening steps speeds up the discovery process. Classical approaches used for virtual screening are not always efficient, being affected, for example, by the use of specific force field [25, 14] or the quality of structural data available, thus affecting the whole discovery process from hit identification to hit optimisation steps towards the lead compound. Hit discovery is the very first step of a drug discovery workflow, and chemists use it as the starting point to optimise molecules rationally up to the pre-clinical candidate. The molecular optimisation steps involves different techniques such as hit expansion, (bio)isosteric replacement or a combination of them [11]. Hit expansion, for instance, is one of the most widely used techniques for optimisation and it is based on the synthesis of analogues to the selected small molecule by replacing chemical patterns, guided by data obtained from Structure Activity Relationship (SAR) analysis [2]. (Bio)isosteric replacement is a technique based on the definition of isosterism introduced by Langmuir in 1919, during his studies of the physicochemical properties of atoms and chemical groups. He noted that atoms with similar properties trigger the same biological effect; for example, compounds with the same number of atoms and electrons (e.g. N2, CO, N2O, etc.) are isosteric and trigger the same biological behaviour. On the basis of the same electron arrays, the chemical groups have been divided into 21 groups that are used to optimise bioactive molecules by increasing their efficiency as [8] drugs. Clearly, optimisation requires an understanding of the relevant chemical features within the small molecule. This step is usually handled by the computational chemist, who recognizes and consciously acts to improve the molecule’s activity in accordance with their knowledge and expertise in the field.
@@
In this context, Graph Neural Networks (GNN) [22, 13] have emerged as powerful tools in Drug Discovery and Virtual Screening due to their ability to model complex molecular structures and interactions. By leveraging the graph-based representation of molecules to capture local and global structural information, GNNs offer a promising approach for predicting molecular properties, bioactivity and drug-target interactions or guide the optimisation steps described above. [10, 34, 28, 30, 24, 5, 27].
Even if the graph representation of a molecule is self-explainable, the precise relation between the relevant molecular properties and the network prediction is not, and it suffers from the same issues observed in classical Convolutional Neural Networks devoted to images. As a consequence, graph oriented XAI techniques have emerged in the Drug Discovery domain. A very recent work with a good literary review is the one by Proietti et al. [18] that addresses the Concept Withening (CW) technique to investigate the decision process of a GCNN aimed at discovering relevant molecular properties on a plethora of benchmark data sets. CW has been presented in [6], and it is a representation learning technique aimed at aligning the axes in the latent space close to some known “concepts” (i.e. group of features) in the data space. Proietti and colleagues use GNNExplainer [31] to learn the contributions of the concepts in the QSAR model predictions. GNNExplainer is an explainer capable of identifying compact subgraphs and small subsets of features through the use of Mutual Information (MI) allowing the GNNExplainer to be formulated as:
(1) $\max_{G_{S}}MI\left(Y,\left(G_{S},X_{S}\right)\right)=H(Y)-H\left(Y\mid G=G_{S},X=X_{S}\right)$
The MI term is computed between the predicted label distribution $Y$ and the combination of subgraph $G_{S}$ and selected features $X_{S}$. This term gauges the dependence between the two variables, measuring how much information about $Y$ is gained by knowing both the subgraph $G_{S}$ and the features $X_{S}$. $H_{Y}$ represents the entropy term, quantifying the overall uncertainty in predicting $Y$ without considering subgraphs or features. The last term, $H\left(Y\mid G=G_{S},X=X_{S}\right)$ denotes the conditional entropy of $Y$ given the subgraph $G_{S}$ and the features $X_{S}$. By maximing the mutual information $MI$ the model seeks the subgraph $G_{S}$ and the features subset $X_{S}$ that provide the most information about $Y$. This is achieved by minimizing the conditional entropy $H\left(Y\mid G=G_{S},X=X_{S}\right)$, indicating that knowing the chosen subgraph and features reduces uncertainty in predicting $Y$ as much as possible.
The SHapley Additive exPlanation (SHAP) is another well known XAI framework that is used also in the Drug Discovery domain. In this context, SHAP aims to explain the contribution of individual nodes and edges (features) in a GNN to the final prediction for a molecule. To assess a node or edge’s importance, SHAP creates a coalition centered around that specific feature. This coalition includes the node/edge itself and all its descendant nodes and edges in the graph. SHAP then compares the model’s prediction for the molecule with and without the entire coalition. This difference reflects the impact of the target node/edge and its connected features on the prediction. By analyzing SHAP values for different coalitions centered on individual nodes and edges, researchers gain insights into the relative importance of each feature for the GNN’s prediction. In the equation 2 $g$ is the explanation model, $z^{\prime}\in{0,1}^{M}$ is the coalition vector, $M$ is the maximum coalition size and $\phi_{j}\in\mathbb{R}$ is the $j$-th feature attribute.
(2) $g\left(z^{\prime}\right)=\phi_{0}+\sum_{j=1}^{M}\phi_{j}z_{j}^{\prime}$
The equation shows how the SHAP framework breaks down model predictions into contributions from single features or feature combinations, giving each a weight determined by its relative importance.
SHAP is utilized in [20, 19] and relies on SHAP values to exclusively identify the importance of different features, but did not specify their pharmacophoric nature. All the explainers described above, while providing additional information for sub-graphs, do not attain the level of detail provided by our HGE approach.
In the present work, we address the explainability issue as follows. We present at first a Graph Convolutional Neural Network (GCNN), which prioritise molecules towards a set of 20 kinases with a very high rate of precision, and also supports the computational chemist in the selection of the most important chemical features in the host-guest recognition. The proposed neural architecture provides a very good bioactivity classification (F1 score $>$ 98% on average).
The GCNN outlined above has been developed as a typical VS application, and it is the test architecture for a suitable Hierarchical Graph Explainability (HGE) procedure, which is the main contribution of this work, and it is inspired to the well known Gradient-weighted Class Activation Mapping (Grad-CAM) [23]. Grad-CAM is an explainer used in image classification through Convolutional Neural Networks (CNN) and it generates a heat-map for each class to highlight the parts of an image that contribute to the recognition of that class. At each layer where Grad-CAM is applied, the role of each feature map in explaining the class is weighted using the gradients backpropagated to the layer itself. In [17] GCNN-Explainability, a graph based version of Grad-CAM is proposed, that uses a Grad-CAM layer at the end of the GCNN architecture to identify the substructures that are relevant to the individual classification and thus characteristic of a given class. The proposed HGE approach, extends this explainer by deepening the analysis of the relevant sub-graphs that are directly related to the chemical functionality of the identified moieties. Our molecular explainer provides a graphical representation highlighting the atoms and/or chemical groups with a key role in the discriminative event. We used four macro-groups of pharmacophoric features (H-bond donor or acceptor, hydrophobic, aromatic) [7] in this respect.
@@
As already mentioned above, Grad-CAM is applied only to the last layer of the classifier whose feature maps convey aggregated position-independent information on the features extracted by the deep layers of the network. Rather than providing a global explanation of the classifier behaviour, the key idea behind our approach is trying to explain how the different molecular moieties are involved in predicting a high degree bioactivity. These structures are expected to play a crucial role in binding to the target. As a consequence, we used a hierarchy of explainers placed at different layers to catch information directly at atom, ring, and whole molecule level by leveraging the message passing mechanism that is used in GNNs. The explainer hierarchy for a target is applied purposely to a binary version of our classifier. Such a network has a similar layer arrangement as the 20-target network, but the layers’ sizes have been designed ad hoc, and a novel training on the selected target has been performed to achieve high accuracy, while avoiding the inherent bias of a multi-target prediction with respect to the different classes. In this paper we present the implementation of the whole explainability procedure for CDK1, which is a well-known regulator of the cell cycle, and has a crucial role in tumor genesis. Our explainability results for nineteen well known drugs that are active on CDK1 where visually inspected and compared to literature results or molecular docking results when data from literature were not available. Docking protocols were assessed through cognate docking experiments and results were validated in comparison with data available from the literature ( known binding mode and crucial protein residues for interaction). The proposed approach is exactly the same for each target so this choice does not reduce the generality of our results.
@@ 2.1 Data preparation
The set of atom features was later augmented with the 3D coordinates of each atom in active configuration, for the sake of training the single-class architecture used for the hierarchical explainability analysis. These coordinates were calculated minimizing their conformation and protonaiting molecules using Ligprep [1] with $pH=7.4$. Edges information concerns the type of bond (double or triple bond) and whether the bond is within a cycle. The size of each molecular graph had several nodes and arcs which vary according to the size of the molecule.
Canonical SMILES were recovered from the CHEMBL24.1 database [33] searching for all the small molecules whose activity on the target protein is referenced by IC50, Ki, and Kd values. We used a fixed threshold activity value on each target, in order to generate the molecular graphs through the use of the RDKit library, which allows obtaining all the features previously described. The final dataset consists of 89373 small molecules and has a 1:100 active/inactive ratio for the smallest class that is CLK2. This ratio allows profound assessing of the model generalization abilities as it has more or less the same size of a real world screening task. Training, validation, and test sets were created using a 80%:10%:10% ratio for each target, in order to minimize the comparison bias.
@@ 2.2 Graph Convolutional Neural Network
The single target architecture differs significantly from the first one as regards both the number and the size of each layer. It is well known that the global efficacy of a model for Virtual Screening is measured through the ability of correctly prioritizing the True Positives, so the main aim of the network is providing a very high sensitivity with respect to the inhibitors of the target that is CDK1 in our experiments. In turn, stressing the network towards high sensitivity performance, makes it able to provide precise explanations about the molecular moieties involved in the inhibition. To achieve this result we operated in two ways: enrichment of the information encoded at each node, and accurate design of the model capacity through a change in the dimensions of the layers. The 3-D coordinates of each atom were computed starting from the minimized conformation of the molecule, and they were added to the node features. Moreover, we selected a tighter network than the multi-target one because the task is only a binary classification.
The GC layers in the single target network elaborate the features in a U-Net [21] fashion, first enlarging, and then tightening the channels at each layer until a final 32 features representation is obtained. Seven GC blocks are used where the channels of each GC layer are arranged as follows: 32, 64, 128, 256, 128, 64, and again 32. Layer normalization and ReLU are obviously used in each block. After the GAP, four fully connected layers where used whose channels are respectively 32, 32, 16, and finally 1 that is fed to a sigmoid to estimate the probability for the input to be active on the target. The whole architecture of the single target network is reported in Figure 2 along with an example of the output of the explainers used in the proposed hierarchy.
@@ 2.3 Explainability analysis
Bioactivity prediction is a classification task where molecules are labelled as positive or negative if they trigger “globally” a particular biological event or not. Many factors are involved in bioactivity apart from the presence of the pharmacophoric moieties, molecular chirality, solubility, and so on. All of them can be captured by training the GCNN on a three-dimensional molecular representation with rich information at the nodes, but can not be directly related to just “positive” or “negative” Grad-CAM values at each node. As a consequence we removed ReLU for our explainability task. Moreover, pharmacophoric information is strictly related to atom connectivity within the molecular structure. In view of this consideration we evaluated Grad-CAM multiple times at different depths in the convolutional segment of the network. We used the classical Grad-CAM arrangement after the last layer of the network that is the 7-th one. This explainer is aimed at highlighting relevant features globally, and it is not able to provide information about subtle moieties like acceptors or donors. To reach this aim, we placed also a Grad-CAM explainer at the very first GC layer to consider the atoms’ impact in a 1-neighbourhood that is at a bond level. Another explainer was placed after the third layer to catch information related to 3-neighborhoods. At this level, as we analyze the 3-neighborhood of each atom, the action of an entire substructure like aromatic ring or carbonyl ring can be detected.
@@
The resulting $L_{Grad-CAM}$ distribution has been min-max scaled, and only contributions higher than $0.7$ have been regarded as relevant. The RDKithttps://www.rdkit.org/ software package has been used to devise pharmacophoric labels for each molecular structure, and the final map reports only relevant contributions colored accordingly to the labels. In particular, H-bond acceptor or donor groups are highlighted in yellow, hydrophobic groups are colored in blue, and the aromatic groups are green. Lastly, atoms without an explicit pharmacophoric label where $L_{Grad-CAM}>0.7$ are labelled as “relevant”, and are colored in red. In Section 3.3 a detailed analysis is reported about the structures highlighted by our hierarchical Grad-CAM explainer.
@@ 3.1 Multi-class multi-target classification
Results for the 20-target classifier are shown in Table 1 and Table 2. The first table provides the performance metrics of our network as a classifier. The Balanced Accuracy assesses the network’s ability to make correct predictions despite the class imbalance, while the Sensitivity (also named Recall) focuses solely on the network’s performance in correctly classifying active molecules. The F1-score quantifies how well the model can correctly identify positive instances (Recall) while avoiding false positives (Precision). On the other hand, the Area Under the Curve (AUC) provides a global measure of the model’s ability to discriminate between positive and negative instances, and it is computed from the curve that plots the Sensitivity vs the False Positive Rate varying the threshold to discriminate positives and negatives.
@@
In our results, both the accuracy and sensitivity values indicate that our multi-target network identifies correctly both the positive and negative samples. Particularly, sensitivity increased significantly with respect to the EMBER classifier where it ranged between 0.22 and 0.84. Now, sensitivity has an average value of 0.904, showing that it is much easier for the network to classify positive samples. The comparison between the two networks is summarised in Figure 3 showing the sensitivity for all targets. It is notable that there was a huge improvement in the performance related to sensitivity. Particularly, sensitivity on the CLK2 target raised from 0.22 to 0.938, as it is shown also in Table 1.
Figure 3: Sensitivity comparison for all the targets between the EMBER and our GC multi-target classifier.
F1-Score values show that the network is able to overcome the dataset imbalance while simultaneously reducing the number of false negatives and false positives. The use of a properly balanced loss made by a weighted cross entropy and a soft F-score worked fine for the classification task. Finally, also the values for TP/P and EF (Table 2) are similar to the ones reported for the EMBER classifier.